OpenStudy (anonymous):

Solve by the multiplication with add and subtract method 3x - 2y = 12 2x + y = 1

OpenStudy (anonymous):

first multipy the second equation by 2 [2x+y=1]*2 4x+2y=2 then add the resulted equation with the first equation 3x-2y=12 + 4x+2y=2 --------- 7x=14 divide by 7 x=14/7 x=2 plug in the value of x=2 in one of the equations 2(2)+y=1 4+y=1 y=1-4 y=-3